An electric dipole that has a positive charge of 4.80 × 10^–19 C is separated from a negative charge of the same magnitude by 6.40 × 10^–10 m. The magnitude of the electric field at the midpoint of the dipole is
Answer #1

Electric field at the midpoint of dipole= E=2kq/r^2 here r=3.2*10^-10 m (midpoint)

= 2*9*10^9*4.8*10^-19/(3.2*10^-10)^2
=8.4375*10^-10 N/C